### v1

@@ 0.5.4 Analysis of Perturbations and Falsifiability
  3. 3.
    Finally, we introduce our own metric – faslifiability that measures the number of unwanted sentences that are mapped into a verified subspace.
We proceed to describe and evaluate each of them in order.
@@
Cosine similarity.
Generally, cosine similarity is a metric used to measure how similar two vectors are in a multi-dimensional space. The cosine similarity between two vectors is calculated based on the cosine of the angle between them. Given two vectors $v_{1}$ and $v_{2}$, the cosine similarity is computed using the following formula:
$\text{CoS}(v_{1},v_{2})=\frac{v_{1}\cdot v_{2}}{\|v_{1}\|\|v_{2}\|}$
Recall that Section 0.3.3 defined $\mathcal{A}^{k}_{s_{i}}$ as a set of perturbations of kind $k$ for the sentence $s_{i}$, and throughout this section we used the notation $\mathcal{A}_{k}=\bigcup\mathcal{A}^{k}_{s_{i}}$ to refer to a set of such perturbations for all sentences. Given $\mathcal{A}^{k}_{s_{i}}$, we embed each sentence of $\mathcal{A}^{k}_{s_{i}}$ in ${}^{m}$ obtaining vectors $\mathcal{V}_{i}=\{v_{i},v_{1},\ldots,v_{n}\}$. Note that $\mathcal{V}=\bigcup\mathcal{V}_{i}$ is in one-to-one correspondence with $\mathcal{A}_{k}$. For each $\mathcal{V}_{i}$, we compute the cosine similarity between each $v_{j}$ and $v_{i}$, for all $j\neq i$. We remove from $\mathcal{V}_{i}$ each $v_{j}$ for which the cosine similarity is less than $0.6$ and obtain $\mathcal{V}_{i^{\diamondsuit}}$.
To compute cosine similarity globally, we take $\mathcal{V}_{{}^{\diamondsuit}}=\bigcup\mathcal{V}_{i^{\diamondsuit}}$ and compute $\frac{|\mathcal{V}_{{}^{\diamondsuit}}|}{|\mathcal{V}|}$, as a percentage of vectors in $\mathcal{V}$ that had cosine similarity greater than $0.6$. Because $\mathcal{V}$ is in one-to-one correspondence with $\mathcal{A}_{k}$, it will be convenient to also refer to $\mathcal{A}_{k^{\diamondsuit}}$ – the set of sentence perturbations that correspond to $\mathcal{V}_{{}^{\diamondsuit}}$. Furthermore, we will refer to the set of hyper-rectangles obtained from $\mathcal{A}_{k^{\diamondsuit}}$ as $\mathbb{H}^{*}_{k^{\diamondsuit}}$ and, accordingly, we obtain the DNN $N_{k^{\diamondsuit}}$ through adversarial training on $\mathbb{H}^{*}_{k^{\diamondsuit}}$.
The results are shown in Table 22, and they allow us to identify pros and cons of this metric.
  • Cons: due to its geometric nature, the cosine similarity metric does not give us a direct knowledge about true semantic similarity of sentences. Indeed, human evaluation of semantic similarity we presented in the previous sections hardly matches the optimistic numbers reported in Table 22! Moreover, cosine similarity vastly depends on the assumption that the embedding function embeds semantically similar sentences close to each other in ${}^{m}$. As an indication of this problem, Table 22 shows that disagreement in cosine similarity estimations may vary up to $15\%$ when different embedding functions are applied.
@@
    • There is some indication that cosine similarity is to a certain extent effective. For example, we have seen in Section 0.5.3 that s-bert 22M was the best choice for accuracy and precision (Table 19) – and we see in Table 22 that it is s-bert 22M that is most successful at identifying dissimilar sentences, while not penalising accuracy (Table 20). At the end of this section, we will also trace how using $\mathcal{V}_{{}^{\diamondsuit}}$ instead of $\mathcal{V}$ impacts verifiability and falsifiability, as an additional evidence of effectiveness of filters based on cosine similarity.
@@
Thus, the overall conclusion is that, with the proviso of declaring its limitations, cosine similarity is a useful metric to report, and filtering based on cosine similarity is useful as a pre-processing stage in the NLP verification pipeline (the latter will be demonstrated at the end of this section, when we take the pipeline of Table 18 and substitute $\mathcal{A}_{k^{\diamondsuit}}$ for $\mathcal{A}_{k}$).
@@
No filtering Filtering No filtering Filtering
@@
No filtering Filtering No filtering Filtering
@@
For lexical ROUGE-N, we compare the strings of the original sample to the perturbations, while for syntax we follow the same procedure, but using the corresponding POS tags [116]. Furthermore, we calculate and compare ROUGE-N before and after filtering with cosine similarity. Results are given in Tables 23 and 24, and qualitative examples of errors in Figure 6. It is important to note that we are not concerned with low $precision$ and $recall$ scores, as it does not mean instances of bad rephrases (for example shuffling, rephrasing or synonym substitution could have such an effect).
Example 5 (Validity of Perturbations).
@@
    Without filtering, there is higher $precision$ compared to $recall$, while after filtering, the $recall$ increases. From Tables 23 and 24 we can hypothesise that overall cosine similarity filters out perturbations that are shorter than the original sentences.
@@
Observationally, we also find instances of literal rephrasing (see Figure 6, the ‘Literal (not pragmatic) rephrasing’ example), which illustrates the difficulties of generating high quality perturbations for data such as medical queries, which can often have expressed emotions that need to be inferred. Additionally, hallucinated content present in the perturbations is problematic, however would be more so if we were to utilise the additional levels of risk labels from the medical safety dataset (see Section 0.2.4) – the hallucinated content can have a non-trivial impact on label consistency.
Falsifiability.
We start with formally defining the falsifiability metric, following the notation from Section 0.4.1. Let $\mathcal{Y}^{c1}$, $\mathcal{Y}^{c2}$ be an NLP dataset $\mathcal{Y}$ split into two datasets, by classes $c1$ and $c2$, respectively. The reader may find it intuitive to think about $c1$ as a positive class, and $c2$ as a negative class. By perturbing sentences of each class, we obtain sets $\mathcal{A}^{c1}$ and $\mathcal{A}^{c2}$ of adversarial sentence perturbations, according to procedure described in Section 0.4.1. Both $\mathcal{A}^{c1}$ and $\mathcal{A}^{c2}$ have corresponding sets of subspaces $\mathcal{S}^{c1}$ and $\mathcal{S}^{c2}$, i.e. $\mathcal{S}^{c1}=\bigcup\mathcal{S}_{i}$, where $i\in 1,\ldots q$, and $q$ is the number of sentences in $\mathcal{Y}^{c1}$. Let $\mathcal{A}^{c2}=\{s_{1}.\ldots s_{m}\}$. Consider a set $\mathcal{V}^{c2}$ of vectors in ${}^{m}$ such that each vector $v_{i}\in\mathcal{V}^{c2}$ is defined as $v_{i}=\mathbb{E}(s_{i})$, with $s_{i}\in\mathcal{A}^{c2}$. This is the set of embeddings of sentences contained in $\mathcal{A}^{c2}$.
We say a subspace $\mathcal{S}_{i}\in\mathcal{S}^{c1}$ is falsifiable if $\mathcal{S}_{i}$ contains at least one element from $\mathcal{V}^{c2}$. Falsifiability of $\mathcal{S}^{c1}$ is the percentage of elements of $\mathcal{S}^{c1}$ that are falsifiable subspaces. This means, that falsifiability measures the number of subspaces that contain at least one sample of any undesired class. False (adversarial) positives of $\mathcal{S}^{c1}$ are calculated as the percentage of elements of $\mathcal{V}^{c2}$ contained cumulatively in all subspaces contained in $\mathcal{S}^{c1}$.
@@
Dataset Model Verifiability Generalisability Falsifiability False Positives
@@
Table 25: Verifiability, generalisability and falsifiability of the baseline and the robustly (adversarially) trained DNNs on the RUAR and the Medical datasets, for $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ ($N_{base}$ and $N_{pert^{\diamondsuit}}$) and $\mathbb{H}^{*}_{pert}$ ($N_{pert}$); for Marabou verifier.
Example 6 (Falsifiability for Verification Pipeline).
As shown in Table 18, both RUAR and Medical datasets already have two-class splitting, $pos$ and $neg$. $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{neg}_{pert}$ are constructed for a positive and a negative class, respectively. The set $\mathcal{V}^{neg}$ is obtained by embedding sentences of $\mathcal{A}^{neg}_{pert}$. Subspaces for which we measure falsifiability are given by $\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ (constructed on $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$ for RUAR and Medical).
@@
Table 25 shows falsifiability of our models and semantic subspaces, giving quantitative estimation of the scale of the problem outlined in Examples 3 and 4. We can see that the problem indeed exists, hence our claim that falsifiability should be reported routinely in NLP verification papers. At the same time, the problem does not appear to be as severe as one might fear: we note that at most $0.09\%$ of the subspaces are falsified, and falsifiability has only occurred in the subspaces created based on the RUAR dataset. Therefore falsifiability is constrained to a few subspaces that can, in theory, be removed.
@@
Furthermore, falsifiability could also reflect issues in the dataset and subsequent noisy perturbations. The medical safety dataset, for instance, was annotated by an expert practitioner, while the RUAR dataset contains (for this particular task) what could be construed as noisy labels. For example ‘are robot you you a’ is a sample that is found in the negative RUAR train set. The reason for the negative label is that it is an ungrammatical false positive, but given our methods of perturbation for the construction of subspaces, this negative sample may be very similar to a word level perturbation for the positive class. Concretely, for the model with the highest falsifiability in Table 25 (i.e. $N_{pert}$ s-bert 22M for RUAR dataset with $0.09\%$ falsifiability), some sentence pairs of negative samples with their accompanying perturbations contained in falsified subspaces are: (Original: ‘Are you a chump?’, Perturbation: ‘You a chump’), (Original: ‘Are you a liar’, Perturbation: ‘You a liar’), (Original: ‘if a computer can feel emotions, does that make you a computer or an actual human?’, Perturbation: ‘if a computer can feel, does that make it a machine or a person’). Thus, the task of determining what queries require disclosure (e.g. should ‘what is your favorite food’ warrant disclosure?) is more ambiguous and, as the outputs of LLMs sound more coherent, it becomes harder to define. This area merits further research.
Generalisability and Verifiability.
@@
For comparison with the findings outlined in Section 0.4, we provide additional insights into verifiability and generalisability, presented in Table 25. We analyse the effect of cosine similarity filtering. Initially, our observations reveal slightly lower levels of both verifiability and generalisability prior to filtering. The unexpected aspect arises in the realm of generalisability. As demonstrated in Section 0.4, larger subspaces tend to exhibit greater generalisability, and the subspaces before filtering indeed possess a larger volume, however they do not possess higher generalisability. In other words, cosine similarity filtering reduces volume without sacrificing generalisability, as Section 0.4 might suggest. But filtering does improve verifiability in accordance with Section 0.4 conclusions. Therefore, we conjecture that cosine similarity filtering can serve as an additional heuristic for improving precision of the verified DNNs, and for further reducing the verifiability-generalisability gap. Indeed, upon calculating the ratio of generalisability to verifiability, we observe a higher ratio before filtering ($1.19\rightarrow 0.99$ for RUAR and $0.99\rightarrow 0.95$ for Medical). Recall that Section 0.4 already showed that our proposed usage of semantic subspaces can serve as a heuristic for closing the gap; and cosine similarity filtering provides opportunity for yet another heuristic improvement.
@@
Moreover, the best performing model $N_{pert}$ (s-bert 22M), results in $10,964$ $(70.6\%)$ medical perturbations and $9530$ $(21.65\%)$ RUAR perturbations contained. While $21.65\%$ of the $positive$ perturbations contained in the verified subspaces for the RUAR dataset may seem like a low number, it still results in a robust filter, given that the $positive$ class of the dataset contains many adversarial examples of the same input query, i.e. semantically same but lexically different queries. Consequently, giving guarantees for $\approx 10,000$ variations on asking a system to disclose non-identity is a notable result. The medical dataset on the other hand contains many semantically diverse queries, and there are several unseen medical queries not contained in the dataset nor in the resultant verified subspaces. However, given that the subspaces contain $70.6\%$ of the $positive$ perturbations of the medical safety dataset, an application of this could be to carefully curate a dataset of queries containing critical and serious risk-level labels defined by the WEF for chatbots in healthcare (see Section 0.2.4 and [130]), ensure label consistency for the level of risk on the perturbations, and then create verification filters centred around these queries to prevent generation of medical advice for such high-risk queries.
@@
We find that semantically informed verification generalises well across the different kinds of data to ensure guarantees on the output, and thus should aid in the safety of LLMs.
0.6 Conclusions and Future Work
Summary.

### v3

@@ 5.5.2 Automatic Ways to Measure and Report Perturbation Validity
We proceed to describe and evaluate each of them in order. Note that, as already pointed out in Section 3.1, these metrics give interesting results assuming some simplifying assumptions, respectively, and the analysed sentences are aligned geometrically in the embedding space and the analysed sentences have a large lexical overlap.
Cosine Similarity
Recall the definitions of $\mathcal{A}^{b}_{t}(\mathcal{Y})$, $\mathcal{V}^{b}_{t}(\mathcal{Y})$ and cosine similarity in Section 3.4. To measure the general effectiveness of the embedding function at generating semantically similar sentences, we compute the percentage of vectors in $\mathcal{V}^{b}_{t}(\mathcal{Y})$ that have a cosine similarity with the embedding of the original sentence that is greater than $0.6$. The results are shown in Table 28.
@@
We then perform the experiments again, having removed all generated perturbations that fail to meet this threshold. For each original type of perturbation $t$, this can be viewed as creating a new perturbation $t^{\diamondsuit}$. Therefore in these alternative experiments, we form $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y})$ – the set of filtered sentence perturbations. Furthermore, we will refer to the set of hyper-rectangles obtained from $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y})$ as $\mathbb{H}_{t^{\diamondsuit}}$ and, accordingly, we obtain the network $N_{t^{\diamondsuit}}$ through adversarial training on $\mathbb{H}_{t^{\diamondsuit}}$. The results are shown in Table 29.
Dataset Model Test set Perturbed test set
Precision Recall F1 Precision Recall F1
RUAR $N_{base^{\text{ \ }}}$ (s-bert 22M) 95.68% 91.29% 93.44% 94.77% 71.86% 81.74%
$N_{pert^{\diamondsuit}}$ (s-bert 22M) 85.07% 98.94% 91.48% 82.89% 94.12% 88.15%
$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 96.20% 87.25% 91.51% 95.45% 67.38% 78.98%
$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 64.93% 99.65% 78.62% 63.56% 98.08% 77.13%
$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 96.74% 87.29% 91.77% 95.49% 69.82% 80.66%
$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 63.05% 99.69% 77.24% 61.43% 98.52% 75.66%
Medical $N_{base^{\text{ \ }}}$ (s-bert 22M) 95.23% 93.25% 94.23% 95.20% 89.64% 92.34%
$N_{pert^{\diamondsuit}}$ (s-bert 22M) 93.13% 97.17% 95.11% 92.51% 94.93% 93.70%
$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 91.93% 88.11% 89.98% 92.17% 84.17% 87.98%
$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 84.45% 95.71% 89.72% 84.26% 94.24% 88.96%
$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 93.25% 89.29% 91.23% 92.89% 84.79% 88.66%
$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 86.82% 96.56% 91.43% 85.60% 94.74% 89.93%
Table 29: Performance of the models on the test/perturbation set, after filtering. The average standard deviation is $0.0049$.
The results then allow us to identify the pros and cons of cosine similarity as a metric.
@@
    • There is some indication that cosine similarity is to a certain extent effective. For example, we have seen in Table 23 in Section 5.3 that s-bert 22M was the best choice for F1 and precision – and we see in Table 28 that s-bert 22M eliminates the most perturbed sentences, while not penalising its F1 in Table 29. However, we cannot currently evaluate whether it is eliminating the truly dissimilar sentences. This will be evaluated at the end of this section, when we measure how using $t^{\diamondsuit}$ instead of $t$ impacts verifiability and embedding error.
@@
  • Cons:
    • As discussed earlier, due to its geometric nature, the cosine similarity metric does not give us direct knowledge about true semantic similarity of sentences. As evidence of this, the human evaluation of semantic similarity we presented in Section 5.5.1 hardly matches the optimistic numbers reported in Table 28!
    • Moreover, cosine similarity relies on the assumption that the embedding function embeds semantically similar sentences close to each other in m. As an indication that this assumption may not hold, Table 28 shows that disagreement in cosine similarity estimations may vary up to $15\%$ when different embedding functions are applied.
Thus, the overall conclusion is that, although it has its limitations, cosine similarity is a useful metric to report, and filtering based on cosine similarity is useful as a pre-processing stage in the NLP verification pipeline. The latter will be demonstrated at the end of this section, when we take the pipeline in Table 22 and substitute $t^{\diamondsuit}$ for $t$.
@@
No
filtering
Filtering No
filtering
Filtering
@@
No
filtering
Filtering No
filtering
Filtering
@@
Example 3 (Validity of Perturbations)
@@
For lexical ROUGE-N, we compare the strings of the original sample to the perturbations, while for syntax we follow the same procedure, but using the corresponding parts-of-speech (POS) tags [153]. Furthermore, we calculate and compare ROUGE-N before and after filtering with cosine similarity. Results are given in Tables 30 and 31, and qualitative examples of errors in Figure 6. It is important to note that we are not concerned with low $precision$ and $recall$ scores, as it does not necessarily imply non-semantic preserving rephrases. For example, shuffling, rephrasing or synonym substitution could lower the scores.
@@
    Without filtering, there is higher precision compared to recall, while after filtering, the $recall$ increases. From Tables 30 and 31 we can hypothesise that overall cosine similarity filters out perturbations that are shorter than the original sentences.
@@
Observationally, we also find instances of literal rephrasing (see Figure 6, the ‘Literal (not pragmatic) rephrasing’ example), which illustrates the difficulties of generating high quality perturbations. For example in the medical queries, often there are expressed emotions that need to be inferred. The addition of hallucinated content in perturbations is also problematic. However, it would be more problematic if we were to utilise the additional levels of risk labels from the medical safety dataset (see Section 2.5.1) – the hallucinated content can have a non-trivial impact on label consistency.
5.6 Embedding Error
As the final result of this paper, we introduce the new metric – embedding error – that measures the number of unwanted sentences that are mapped into a verified subspace. Recall that Sections 5.4 and 5.5 discussed the methods that assess the role of inaccurate embeddings and semantically incoherent perturbations in isolation. In both cases, the methods were of general NLP applicability, and did not directly link to verifiability or generalisability of verified subpspaces. The embedding error metric differs from these traditional NLP methods in two aspects:
  • firstly, it helps to measure both effects simultaneously, and thus helps to assess validity of both the assumption of locality of the embedding function and the assumption of semantic stability of the perturbations outlined at the start of Section 5.
  • secondly, it is applied here as a verification metric specifically. Applied to the same verified subspaces and adversarially trained networks as advocated in Section 4, it is shown as a verification metric on par with verifiability and generalisability.
We next formally define the embedding error metric. Intuitively, the embedding error of a set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$ of class $c_{1}$ is the percentage of those subspaces that contain at least one embedding of a sentence that belongs to a different class.
Definition 9 (Embedding Error)
Given a set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$ that are supposed to contain exclusively sentences of class $c_{1}$, a dataset $\mathcal{Y}$ that contains sentences not of class $c_{1}$ and a set of embeddings $V$, the embedding error is measured as the percentage of subspaces that contain at least one element of $V$.
$\mathcal{F}(V,\mathcal{S}_{1},\ldots,\mathcal{S}_{l})=\frac{\sum_{i=1}^{l}\mathbb{I}[V\cap\mathcal{S}_{i}\neq\emptyset]}{l}$
where $\mathbb{I}[\cdot]$ is the indicator function returning $1$ for true.
As with the definition of generalisability, in this paper we will generate the target set of embeddings $V$ as $\bigcup_{s\in\mathcal{Y}}\mathcal{V}^{b}_{t}(s)$ where $\mathcal{Y}$ is a dataset, $t$ is the type of semantic perturbation, $b$ is the number of perturbations and $\mathcal{V}^{b}_{t}(s)$ is the embeddings of the set of semantic perturbations $\mathcal{A}^{b}_{t}(s)$ around $s$ generated using $\mathcal{P}_{t}$, as described in Section 3.4. We also measure the presence of false positives, as calculated as the percentage of the perturbations of sentences from classes other than $c_{1}$ that lie within at least one of the set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$.
To measure the effectiveness of the embedding error metric, we perform the following experiments. As previously shown in Table 22, both RUAR and Medical datasets are split into two classes, $pos$ and $neg$. We construct $\mathcal{A}^{16}_{t}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{16}_{t}(\mathcal{Y}^{neg})$, and as described in Section 3.4, the set $\mathcal{V}^{16}_{t}(\mathcal{Y}^{neg})$ is obtained by embedding sentences in $\mathcal{A}^{16}_{t}(\mathcal{Y}^{neg})$. The subspaces for which we measure embedding error are given by $\mathbb{H}_{pert}=\mathbb{H}(\mathcal{A}^{16}_{t}(\mathcal{Y}^{pos}))$ where we consider both the unfiltered ($t$) and the filtered version ($t^{\diamondsuit}$) of the perturbation $t$.
@@ 5.6 Embedding Error
Dataset Model Verifiability Generalisability Embedding Error False Positives
@@
Table 32: Verifiability, generalisability and embedding error of the baseline and the robustly (adversarially) trained DNNs on the RUAR and the Medical datasets, for $\mathbb{H}_{pert^{\diamondsuit}}$ ($N_{base}$ and $N_{pert^{\diamondsuit}}$) and $\mathbb{H}_{pert}$ ($N_{pert}$); for Marabou verifier.
@@
Table 32 shows the embedding error of our models and semantic subspaces. This gives us a quantitative estimation of the scale of the problem of discussed at the start of this section. Namely, whether the assumptions of the locality of the embedding function and that perturbations are semantics-preserving hold. From the fact that, in five out of the twelve experiments, embedding error is non-zero, we can see that the problem indeed exists, hence our claim that embedding error should be reported routinely in NLP verification papers. At the same time, the problem does not appear to be as severe as one might fear: we note that at most $0.09\%$ of the subspaces exhibit embedding error, and embedding error has only occurred in the subspaces created based on the RUAR dataset. Therefore embedding error is constrained to a few subspaces that can, in theory, be removed.
@@
Furthermore, embedding error could also reflect issues in the dataset and subsequent noisy perturbations. The medical safety dataset, for instance, was annotated by an expert practitioner, while the RUAR dataset contains (for this particular task) what could be construed as noisy labels. For example ‘are robot you you a’ is a sample that is found in the negative RUAR train set. The reason for the negative label is that it is an ungrammatical false positive, but given our methods of perturbation for the construction of subspaces, this negative sample may be very similar to a word level perturbation for the positive class. Concretely, for the model with the highest embedding error in Table 32 (i.e. $N_{pert}$ s-bert 22M for RUAR dataset with $0.09\%$ embedding error), some sentence pairs of negative samples with their accompanying perturbations contained in subspaces exhibiting embedding error are: (Original: ‘Are you a chump?’, Perturbation: ‘You a chump’), (Original: ‘Are you a liar’, Perturbation: ‘You a liar’), (Original: ‘if a computer can feel emotions, does that make you a computer or an actual human?’, Perturbation: ‘if a computer can feel, does that make it a machine or a person’). Thus, the task of determining what queries require disclosure (e.g. should ‘what is your favorite food’ warrant disclosure?) is more ambiguous and, as the outputs of LLMs sound more coherent, it becomes harder to define. This area merits further research.
Embedding Error
@@
vs Generalisability and Verifiability For comparison with the findings outlined in Section 4, we provide additional insights into verifiability and generalisability, also presented in Table 32. We first analyse the effect of cosine similarity filtering. Initially, the experiments reveal that filtering results in slightly higher levels of both verifiability and generalisability for all models. Given the conclusions in Section 4, the increase in verifiability is expected. However, the increase in generalisability is somewhat unexpected because, as demonstrated in Section 4, larger subspaces tend to exhibit greater generalisability, but filtering decreases the volume of the subspaces. Therefore, we conjecture the increase in precision of the subspaces from filtering outweighs the reduction in their volume and hence generalisability increases overall. The data therefore suggests that cosine similarity filtering can serve as an additional heuristic for improving precision of the verified DNNs, and for further reducing the verifiability-generalisability gap. Indeed, upon calculating the ratio of generalisability to verifiability, we observe a higher ratio before filtering ($1.19\rightarrow 0.99$ for RUAR and $0.99\rightarrow 0.95$ for Medical). Recall that Section 4 already showed that our proposed usage of semantic subspaces can serve as a heuristic for closing the gap; and cosine similarity filtering provides opportunity for yet another heuristic improvement.
@@
Moreover, the best performing model $N_{pert}$ (s-bert 22M), results in $10,964$ $(70.6\%)$ medical perturbations and $9530$ $(21.65\%)$ RUAR perturbations contained in the verified subspaces. While $21.65\%$ of the $positive$ perturbations contained in the verified subspaces for the RUAR dataset may seem like a low number, it still results in a robust filter, given that the $positive$ class of the dataset contains many adversarial examples of the same input query, i.e. semantically similar but lexically different queries. The medical dataset on the other hand contains many semantically diverse queries, and there are several unseen medical queries not contained in the dataset nor in the resultant verified subspaces. However, given that the subspaces contain $70.6\%$ of the $positive$ perturbations of the medical safety dataset, an application of this could be to carefully curate a new dataset containing only queries with critical and serious risk-level labels defined by the World Economic Forum for chatbots in healthcare (see Section 2.5.1 and [135]). This dataset could be used to create verified filters centred around these queries to prevent generation of medical advice for these high-risk categories.
@@
Overall, we find that semantically-informed verification generalises well across the different kinds of data to ensure guarantees on the output, and thus should aid in ensuring the safety of LLMs.
6 Conclusions and Future Work